OpenStudy (anonymous):

how do you solve : (root3+root2)^2 ?

OpenStudy (anonymous):

a square root squared is just a normal number

OpenStudy (anonymous):

so it would be 3+2=5

OpenStudy (anonymous):

Ah I see, thank you :)!

myininaya (myininaya):

omg wait jim

jimthompson5910 (jim_thompson5910):

oh wait, nvm lol

myininaya (myininaya):

there is a mistake

jimthompson5910 (jim_thompson5910):

i know, just caught it lol

myininaya (myininaya):

you actually made one

myininaya (myininaya):

i cannot believe this

jimthompson5910 (jim_thompson5910):

\[\large (\sqrt{3}+\sqrt{2})^2\] \[\large (\sqrt{3})^2+2\sqrt{3}\sqrt{2}+(\sqrt{2})^2\] \[\large 3+2\sqrt{6}+2\] \[\large 5+2\sqrt{6}\] So \[\large (\sqrt{3}+\sqrt{2})^2=5+2\sqrt{6}\]

myininaya (myininaya):

ok very good! :) lol

jimthompson5910 (jim_thompson5910):

small mistake with radicals = radical mistake...wait a minute....

OpenStudy (anonymous):

:,( my answer was completely and foolishly wrong...

myininaya (myininaya):

lol poor gandalf

jimthompson5910 (jim_thompson5910):

you shall not pass...lol

OpenStudy (anonymous):

lol! I still get to save middle earth though...

jimthompson5910 (jim_thompson5910):

good point...